# MissionHD: Hyperdimensional Refinement of Distribution-Deficient Reasoning Graphs for Video Anomaly Detection

### v3

@@ Abstract
LLM-generated reasoning graphs, referred to as mission-specific graphs (MSGs), are increasingly used for video anomaly detection (VAD) and recognition (VAR). These MSGs are novel artifacts: they often exhibit skewed connectivity and lack large-scale datasets for pre-training, which makes existing graph structure refinement (GSR) methods ineffective. To address this challenge, we propose HDC-constrained Graph Structure Refinement (HDC-GSR), a paradigm that leverages hyperdimensional computing (HDC) to optimize decodable graph representations without relying on structural-distribution learning. Building on this paradigm, we introduce MissionHD, an HDC framework that encodes graphs with constrained graph-neural operations, aligns them directly with downstream task loss, and decodes refined structures. Experiments on VAD/VAR benchmarks demonstrate that MissionHD-refined graphs consistently improve performance, establishing HDC-GSR as an effective pre-processing step for structured reasoning in video anomaly tasks.
@@ 1 Introduction
Detecting anomalies in human activities via artificial vision is crucial for rapid incident response. To reduce annotation cost, weakly supervised video anomaly detection/recognition (VAD/VAR) has become standard anomalyclip ; vad1_cvpr ; rtfm ; vad2_cvpr ; yun2025missiongnn . A growing body of work uses LLM-generated mission-specific graphs (MSGs) to scaffold reasoning for these tasks luo2024rog ; qiu2025step ; li-etal-2025-graphotter ; yun2025missiongnn , providing interpretable structure while avoiding full end-to-end LLM fine-tuning.
@@
Figure 1: The proposed MissionHD framework refines LLM-generated mission-specific graphs (MSGs) through the HDC-GSR paradigm. By leveraging hyperdimensional encoding and decoding, it aligns graph structures with downstream tasks without relying on direct structural-distribution learning.
Most existing pipelines treat the MSG as fixed, implicitly assuming the initial structure is suitable for the downstream video task. In practice, LLM-generated graphs are generic and may include irrelevant nodes/edges; more importantly, their connectivity can be highly skewed (e.g., tens of thousands of input nodes represent video frames feeding a small set of reasoning nodes), so the graph lacks a rich, learnable structural distribution. Furthermore, unlike graphs derived from established domains (e.g., social or citation networks), LLM-generated MSGs are novel artifacts that lack large-scale datasets for pre-training, making conventional GSR methods infeasible. Classical Graph Structure Refinement (GSR) gsr_23 ; ecl-gsr assumes graphs are faithful samples from well-formed distributions and relies on unsupervised/self-supervised signals to “denoise” edges. These assumptions do not hold for LLM-generated MSGs. A major underlying reason is that prior GSR frameworks exhibit limitation on jointly (i) learn a task-aligned graph representation and (ii) decode that representation back to an explicit graph structure, avoiding direct learning of graph distribution; instead, they implicitly depend on the existence of a learnable structural distribution, which is impractical in our setting.
@@
Recent work shows that Hyperdimensional Computing (HDC) provides strong encode–decode capabilities for structured data kleyko2022survey ; poduval2022graphd ; dalvi2025hyperdimensional . Operating in a single high-dimensional space, HDC composes information via simple algebra—binding (element-wise multiplication) and bundling (element-wise addition)—to produce compact, decodable codes that are robust to noise and can operate without gradient-based training. This suggests a path to task-aligned, decodable graph optimization that sidesteps structural-distribution learning entirely: align the representation with task loss, then decode edits back to the graph.
We propose HDC-constrained Graph Structure Refinement (HDC-GSR), which directly optimizes a decodable graph code in a single hyperdimensional space using the downstream task loss. Our encoding uses HDC operations that form a constrained GNN encoder: layer hypervectors act as diagonal linear maps in the shared space (i.e., an MLP with diagonal weights), binding realizes message passing, and bundling realizes aggregation. After aligning the code with the downstream loss, we decode edge contributions via HDC similarity probes and update the MSG—thereby refining structure to match task signals while avoiding impractical structural-distribution modeling.
We instantiate HDC-GSR for video anomaly tasks in MissionHD, which efficiently encodes hierarchical reasoning paths into compositional hypervectors and decodes edge scores for graph editing under weak supervision. Replacing the original LLM-generated MSGs with MissionHD-refined graphs yields consistent gains for standard reasoning models on VAD/VAR.
Our key contributions are as follows:
  • We introduce HDC-constrained Graph Structure Refinement (HDC-GSR) for LLM-generated mission-specific graphs in video anomaly tasks, aligning a decodable graph code with task loss and eliminating the need to learn a structural graph distribution.
  • We highlight that LLM-generated MSGs are novel artifacts that lack large-scale datasets for pre-training, making conventional GSR impractical and motivating our HDC-based approach.
@@
  • We develop a compact encode–decode procedure in a single HDC space: binding/bundling realize a constrained GNN (diagonal layer maps), and decoding scores/edits edges by compositional similarity.
  • We present MissionHD, an HDC-GSR instantiation for VAD/VAR, and show that refining LLM-generated graphs with MissionHDconsistently improves downstream reasoning performance.
@@ 2.1 Hyperdimensional Computing
HDC is a vector-symbolic computational approach inspired by human cognitive processes that encodes concepts as high-dimensional vectors, typically exceeding $1,000$ dimensions kanerva2009hyperdimensional ; thomas2021theoretical . Leveraging HDC-specific mathematical operations—namely bundling and binding—this framework enables a range of machine learning tasks such as classification, regression, and link prediction across diverse modalities, including graphs poduval2022graphd ; chen2024hdreason ; jeong2025exploiting . In this section, we provide a concise overview of the fundamental operations that form the foundation of our proposed method.
@@
In prior HDC approaches for general graph structures, such as GrapHD poduval2022graphd , an unweighted graph $G=(\mathcal{V},\mathcal{E})$ is encoded using a composition of node and neighborhood information. Each node $v_{i}\in\mathcal{V}$ is assigned a random base hypervector $\mathbf{E}_{i}\in\mathcal{H}$. Its local neighborhood is captured by a memory vector $\mathbf{M}_{i}=\bigoplus_{v_{j}\in\mathcal{N}(i)}\mathbf{E}_{j}$, where $\mathcal{N}(i)$ denotes the set of adjacent nodes. The overall graph hypervector is then computed as:
@@ 2.2 Visual Representation Learning Using Graphs
A graph is a powerful data structure for capturing relationships between entities. The application of graphs to solve vision tasks has been widely studied, ranging from obtaining scene graphs to represent connections between objects in a single image xu2017scenegraph ; huang2025building , to organizing both verbal and visual clues for question answering vqa-gnn , and to video anomaly detection zhong2019graph_mil_video ; xdv ; masuakwa2025pv . In particular, novel methods that facilitate the decision-making process of machine-learning models with the aid of knowledge graphs have proven effective across various tasks luo2024rog ; qiu2025step ; li-etal-2025-graphotter . Among these, MissionGNN yun2025missiongnn has shown promising results in video anomaly detection. They construct mission-specific graphs (MSGs) layered directed acyclic graphs (DAGs) $G=(\mathcal{V},\mathcal{E})$ constructed dynamically from task instructions, external knowledge bases. Nodes $v_{k}^{(i)}\in\mathcal{V}$ are assigned to semantic levels $i\in\{1,\dots,\ell\}$, and edges are only allowed between adjacent layers, i.e., $(v_{k}^{(i)},v_{t}^{(i+1)})\in\mathcal{E}$, enforcing a level-monotonic hierarchy. Each node encodes a task-relevant concept such as object attributes, spatial relations, or symbolic goals, and is associated with a feature vector $\mathbf{x}^{(i)}_{k}\in\mathbb{R}^{m}$. Reasoning in MSGs proceeds along reasoning paths—maximal sequences $(v_{k_{1}}^{(1)},\dots,v_{k_{\ell}}^{(\ell)})$ tracing causal chains from sensor-level observations to high-level decisions masuakwa2025pv . For example, a path might encode: “mug on table” $\rightarrow$ “is dirty” $\rightarrow$ “move to sink”.
However, a core limitation in the MSG-based approaches is that their generated MSG nodes and edges reflect what LLMs “see,” rather than what the visual backbones and subsequent decision-making layers of the VAD model observe. Recent research indicates that LLMs can indeed imagine visual worlds  sharma2024vision ; nonetheless, there remains a significant discrepancy between actual visual data and the visual clues inferred by LLMs, as these models are pre-trained predominantly on extensive text corpora. Motivated by this limitation, we propose MissionHD, a novel framework that dynamically reformulates MSGs during each training epoch, aligning them more closely with the representations perceived by the final decision-making layer, aided by HDC.
@@ 3 Methodology
Our framework, MissionHD, refines LLM-generated reasoning graphs by learning their optimal structure for downstream visual tasks. As illustrated in Figure 2, the core of our approach is an end-to-end trainable framework that leverages Hyperdimensional Computing (HDC) to create a task-aligned, decodable representation of the graph. We first encode the graph’s structure and semantics into a single hypervector, which is then optimized via a downstream task loss that implicitly learns structural edits. Finally, we decode this optimized hypervector back into a refined graph structure by scoring and selecting the most salient edges.
@@ 5 Experiments
Adjacency matrix
90.52
-1.62
79.31
+13.72
81.49
+0.02
86.58
-2.87
41.32
+13.9
63.90
+11.6
72.79
+0.12
93.85
+3.37
97.37
+2.26
56.94
+9.3
94.43
+1.55
70.87
+3.29
80.67
-7.93
82.72
+6.14
77.09
+4.19
GCN
93.60
+1.46
65.55
-0.04
79.04
-2.43
84.15
-5.3
51.69
+24.27
77.37
+25.07
85.91
+13.24
90.59
+0.11
97.09
+1.98
65.58
+17.94
92.39
-0.49
71.90
+4.32
91.91
+3.31
85.50
+8.92
79.89
+6.99
GAT
93.81
+1.67
75.27
+9.68
81.95
+0.48
78.99
-10.46
33.20
+5.78
78.52
+26.22
82.19
+9.52
94.16
+3.68
96.15
+1.04
59.23
+11.59
90.83
-2.05
69.47
+1.89
92.86
+4.26
86.45
+9.87
78.40
+5.5
ECL-GSR
95.11
+2.97
71.53
+5.94
87.77
+6.3
78.02
-11.43
66.98
+39.56
82.04
+29.74
79.73
+7.06
95.41
+4.93
95.26
+0.15
64.61
+16.97
92.11
-0.77
58.94
-8.64
95.27
+6.67
81.60
+5.02
80.71
+7.80
MissionHD
95.67
+3.53
88.30
+22.71
84.63
+3.16
89.19
-0.26
47.09
+19.67
83.19
+30.89
78.29
+5.62
91.42
+0.94
97.58
+2.47
62.55
+14.91
93.24
+0.36
72.32
+4.74
92.23
+3.63
93.30
+16.72
82.56
+9.66
@@
Adjacency matrix
92.47
-0.23
89.52
+0.61
98.06
+0.27
94.84
+0.14
91.37
+0.45
94.27
-0.72
91.21
-0.39
93.21
+0.06
GCN
92.14
-0.56
89.51
+0.6
98.88
+1.09
94.21
-0.49
91.00
+0.08
95.50
+0.51
91.87
+0.27
93.49
+0.34
GAT
92.79
+0.09
90.49
+1.58
97.90
+0.11
94.03
-0.67
91.76
+0.84
95.13
+0.14
91.92
+0.32
93.53
+0.38
ECL-GSR
93.32
+0.62
84.75
-4.16
97.91
+0.12
94.60
-0.1
92.59
+1.67
95.96
+0.97
91.88
+0.28
92.94
-0.2
MissionHD
93.41
+0.71
91.07
+2.16
98.32
+0.53
94.45
-0.25
91.29
+0.37
95.16
+0.17
91.95
+0.35
93.71
+0.56
@@ 5.1 Experimental Settings
We evaluate our proposed MissionHD framework on the challenging tasks of VAD and VAR with an NVIDIA A100 GPU. Our experimental setup closely follows the methodology established in the previous paper yun2025missiongnn . We conduct experiments on two widely-used benchmarks for real-world anomaly detection: UCF-Crime sultani2018real and XD-Violence xdv , both of which consist of videos captured from surveillance cameras. For pre-trained joint embedding model, we used ImageBind-Huge girdhar2023imagebind . Also, GPT-4 is used to generate LLM-generated MSGs. Performance is measured using the Area Under the Receiver Operating Characteristic Curve (AUC) for individual classes and the mean AUC (mAUC) across all anomaly classes to provide a comprehensive assessment. Key hyperparameters for our model include a hyperdimensional space of $D=10,000$, a learnable dimensionality of $d=32$ for the learnable hypervectors, a hidden dimensionality of 32 for the final decision transformer layer, and a learning rate of $10^{-5}$. We explore graph refinement thresholds of 0.1, 0.2, 0.4, and 0.8. While our primary evaluation focuses on VAD/VAR, the MissionHD framework is designed to be task-agnostic and can be readily applied to other domains requiring structured, interpretable reasoning over learned knowledge graphs.

### v4

@@ Abstract
LLM-generated reasoning graphs, referred to as mission-specific graphs (MSGs), are increasingly used for video anomaly detection (VAD) and recognition (VAR). However, they are typically treated as fixed despite being generic and distribution-deficient. Conventional graph structure refinement (GSR) methods are ill-suited to this setting, as they rely on learning structural distributions that are absent in LLM-generated graphs. We propose HDC-constrained Graph Structure Refinement (HDC-GSR), a new paradigm that directly optimizes a decodable, task-aligned graph representation in a single hyperdimensional space without distribution modeling. Leveraging Hyperdimensional Computing (HDC), our framework encodes graphs via binding and bundling operations, aligns the resulting graph code with downstream loss, and decodes edge contributions to refine the structure. We instantiate this approach as MissionHD for weakly supervised VAD/VAR and demonstrate consistent performance gains on benchmark datasets.
@@ 1 Introduction
Detecting anomalies in human activities through artificial vision is essential for enabling rapid and reliable incident response. To make such systems practical at scale, researchers have focused on weakly supervised video anomaly detection and recognition (VAD/VAR), which mitigates the high cost of frame-level annotations while maintaining competitive performance Zanella et al. (2024); Yang et al. (2024); Tian et al. (2021); Lv et al. (2023); Yun et al. (2025). Beyond reducing annotation overhead, recent research has also shifted toward mitigating the substantial computational cost of heavy backbone models, particularly large language models (LLMs), which offer strong anomaly detection capabilities but incur significant overhead. A growing body of work addresses this challenge by leveraging LLM-generated mission-specific graphs (MSGs) to scaffold structured reasoning Luo et al. (2024); Qiu et al. (2025); Li et al. (2025); Yun et al. (2025). By distilling high-level task knowledge into compact graph structures, these approaches enable interpretable intermediate representations while avoiding repeated, costly end-to-end LLM inference or fine-tuning.
@@
Figure 1: The proposed MissionHD framework refines LLM-generated MSGs through the HDC-GSR paradigm. By leveraging hyperdimensional encoding and decoding, it aligns graph structures with downstream tasks without relying on direct structural-distribution learning.
Despite their promise, existing approaches typically treat the generated MSG as fixed, adopting it without further adaptation and implicitly assuming that the LLM-produced structure is well aligned with downstream objectives. In practice, such graphs are often generic and may contain irrelevant nodes or edges, which can degrade task performance. This naturally raises the need to refine LLM-generated MSGs so that their structure better supports downstream objectives.
Existing works on Graph Structure Refinement (GSR) Zhao et al. (2023); Zeng et al. (2025a) typically learn a graph structural distribution, either by refining the given graph itself or by leveraging large-scale graph datasets that share the same distribution as the target graphs. However, LLM-generated graphs for downstream tasks are not well-suited to such approaches for two main reasons. First, they lack a meaningful structural distribution. Since LLM-generated graphs are not observed from real-world domains (e.g., social or citation networks), it is unclear whether they exhibit stable or learnable structural regularities. Moreover, LLMs may produce graphs that are difficult to model structurally, such as graphs that are too small or that exhibit highly skewed connectivity (e.g., tens of thousands of input nodes representing video frames feeding into only a small set of reasoning nodes). Second, LLM-generated MSGs are novel artifacts that lack large-scale datasets sharing similar structural distributions. Unlike graphs derived from established domains, there are no large-scale corpora of LLM-generated graphs available for pre-training, rendering conventional GSR methods impractical. To address these limitations—stemming from the difficulty of learning structural distributions and the novel nature of MSGs—we identify a fundamental shortcoming in prior GSR frameworks: they focus on learning graph distributions directly, rather than learning a task-aligned graph representation. In the context of LLM-generated MSGs, where structural distributions are weak or nonexistent, it is more appropriate to learn representations that are directly aligned with downstream task objectives, instead of relying on distribution-level modeling.
@@
Motivated by this, we seek an alternative paradigm that avoids structural-distribution learning and directly optimizes task-aligned graph representations. Recent work shows that Hyperdimensional Computing (HDC) provides strong encode–decode capabilities for structured data Kleyko et al. (2022); Poduval et al. (2022); Dalvi and Honavar (2025). Operating in a single high-dimensional space, HDC composes information through simple algebraic operations—binding (element-wise multiplication) and bundling (element-wise addition)—to produce compact and decodable representations that are robust to noise and do not require modeling an explicit structural distribution. These properties suggest a principled path toward task-aligned, decodable graph optimization: instead of learning a graph distribution, we encode the graph into a hyperdimensional representation, align this representation directly with the downstream task loss, and subsequently decode the optimized representation back into an explicit graph structure. In this way, graph refinement is driven by task supervision rather than by assumptions about an underlying structural distribution.
We formalize this idea as HDC-constrained Graph Structure Refinement (HDC-GSR), a general framework that directly optimizes a decodable graph representation in a single hyperdimensional space under downstream task supervision. In HDC-GSR, HDC operations induce a constrained GNN encoder: layer hypervectors function as diagonal linear maps in the shared space (i.e., an MLP with diagonal weights), binding implements message passing, and bundling performs aggregation. The resulting graph code is aligned with the task loss, and edge contributions are then decoded via HDC similarity probes to update the explicit graph structure. Through this encode–align–decode process, graph refinement is guided directly by task signals, eliminating the need for structural-distribution modeling. Building on this paradigm, we instantiate HDC-GSR for weakly supervised VAD/VAR in MissionHD, which efficiently encodes hierarchical reasoning paths into compositional hypervectors and decodes edge scores for task-driven graph editing. Replacing the original LLM-generated MSGs with MissionHD-refined graphs yields consistent performance gains on video anomaly tasks.
Our key contributions are summarized as follows:
  • We propose HDC-constrained Graph Structure Refinement (HDC-GSR), a new paradigm for refining LLM-generated mission-specific graphs that directly optimizes a decodable, task-aligned graph representation without relying on structural-distribution learning.
@@
  • We design a compact encode–align–decode framework in a single hyperdimensional space, where binding and bundling induce a constrained GNN (with diagonal layer maps), and edge refinement is achieved through compositional similarity-based decoding.
  • We instantiate this paradigm as MissionHD for weakly supervised VAD/VAR and demonstrate consistent improvements over strong baselines on benchmark datasets.
@@ 2.1 Hyperdimensional Computing
HDC is a vector-symbolic computational approach inspired by human cognitive processes that encodes concepts as high-dimensional vectors, typically exceeding $1,000$ dimensions Kanerva (2009); Thomas et al. (2021). Leveraging HDC-specific mathematical operations—namely bundling and binding—this framework enables a range of machine learning tasks such as classification, regression, and link prediction across diverse modalities, including graphs Poduval et al. (2022); Chen et al. (2024); Jeong et al. (2025). In this section, we provide a concise overview of the fundamental operations that form the foundation of our proposed method.
@@
In prior HDC approaches for general graph structures, such as GrapHD Poduval et al. (2022), an unweighted graph $G=(\mathcal{V},\mathcal{E})$ is encoded using a composition of node and neighborhood information. Each node $v_{i}\in\mathcal{V}$ is assigned a random base hypervector $\mathbf{E}_{i}\in\mathcal{H}$. Its local neighborhood is captured by a memory vector $\mathbf{M}_{i}=\bigoplus_{v_{j}\in\mathcal{N}(i)}\mathbf{E}_{j}$, where $\mathcal{N}(i)$ denotes the set of adjacent nodes. The overall graph hypervector is then computed as:
@@ 2.2 Visual Representation Learning Using Graphs
A graph is a powerful data structure for capturing relationships between entities. The application of graphs to solve vision tasks has been widely studied, ranging from obtaining scene graphs to represent connections between objects in a single image Xu et al. (2017); Huang et al. (2025), to organizing both verbal and visual clues for question answering Wang et al. (2023), and to video anomaly detection Zhong et al. (2019); Wu et al. (2020); Masuakwa et al. (2025). In particular, novel methods that facilitate the decision-making process of machine-learning models with the aid of knowledge graphs have proven effective across various tasks Luo et al. (2024); Qiu et al. (2025); Li et al. (2025). Among these, MissionGNN Yun et al. (2025) has shown promising results in video anomaly detection. They construct mission-specific graphs (MSGs) layered directed acyclic graphs (DAGs) $G=(\mathcal{V},\mathcal{E})$ constructed dynamically from task instructions, external knowledge bases. Nodes $v_{k}^{(i)}\in\mathcal{V}$ are assigned to semantic levels $i\in\{1,\dots,\ell\}$, and edges are only allowed between adjacent layers, i.e., $(v_{k}^{(i)},v_{t}^{(i+1)})\in\mathcal{E}$, enforcing a level-monotonic hierarchy. Each node encodes a task-relevant concept such as object attributes, spatial relations, or symbolic goals, and is associated with a feature vector $\mathbf{x}^{(i)}_{k}\in\mathbb{R}^{m}$. Reasoning in MSGs proceeds along reasoning paths—maximal sequences $(v_{k_{1}}^{(1)},\dots,v_{k_{\ell}}^{(\ell)})$ tracing causal chains from sensor-level observations to high-level decisions Masuakwa et al. (2025). For example, a path might encode: “mug on table” $\rightarrow$ “is dirty” $\rightarrow$ “move to sink”.
However, a core limitation in the MSG-based approaches is that their generated MSG nodes and edges reflect what LLMs “see,” rather than what the visual backbones and subsequent decision-making layers of the VAD model observe. Recent research indicates that LLMs can indeed imagine visual worlds  Sharma et al. (2024); nonetheless, there remains a significant discrepancy between actual visual data and the visual clues inferred by LLMs, as these models are pre-trained predominantly on extensive text corpora. Motivated by this limitation, we propose MissionHD, a novel framework that dynamically reformulates MSGs during each training epoch, aligning them more closely with the representations perceived by the final decision-making layer, aided by HDC.
@@ 3 Methodology
Our framework, MissionHD, refines LLM-generated reasoning graphs by learning their optimal structure for downstream visual tasks. As illustrated in Figure 2, the core of our approach is an end-to-end trainable framework that leverages HDC to create a task-aligned, decodable representation of the graph. We first encode the graph’s structure and semantics into a single hypervector, which is then optimized via a downstream task loss that implicitly learns structural edits. Finally, we decode this optimized hypervector back into a refined graph structure by scoring and selecting the most salient edges.
@@ 5 Experiments
Adjacency matrix 90.52 -1.62 79.31 +13.72 81.49 +0.02 86.58 -2.87 41.32 +13.9 63.90 +11.6 72.79 +0.12 93.85 +3.37 97.37 +2.26 56.94 +9.3 94.43 +1.55 70.87 +3.29 80.67 -7.93 82.72 +6.14 77.09 +4.19
GCN 93.60 +1.46 65.55 -0.04 79.04 -2.43 84.15 -5.3 51.69 +24.27 77.37 +25.07 85.91 +13.24 90.59 +0.11 97.09 +1.98 65.58 +17.94 92.39 -0.49 71.90 +4.32 91.91 +3.31 85.50 +8.92 79.89 +6.99
GAT 93.81 +1.67 75.27 +9.68 81.95 +0.48 78.99 -10.46 33.20 +5.78 78.52 +26.22 82.19 +9.52 94.16 +3.68 96.15 +1.04 59.23 +11.59 90.83 -2.05 69.47 +1.89 92.86 +4.26 86.45 +9.87 78.40 +5.5
ECL-GSR 95.11 +2.97 71.53 +5.94 87.77 +6.3 78.02 -11.43 66.98 +39.56 82.04 +29.74 79.73 +7.06 95.41 +4.93 95.26 +0.15 64.61 +16.97 92.11 -0.77 58.94 -8.64 95.27 +6.67 81.60 +5.02 80.71 +7.80
MissionHD 95.67 +3.53 88.30 +22.71 84.63 +3.16 89.19 -0.26 47.09 +19.67 83.19 +30.89 78.29 +5.62 91.42 +0.94 97.58 +2.47 62.55 +14.91 93.24 +0.36 72.32 +4.74 92.23 +3.63 93.30 +16.72 82.56 +9.66
@@
Adjacency matrix 92.47 -0.23 89.52 +0.61 98.06 +0.27 94.84 +0.14 91.37 +0.45 94.27 -0.72 91.21 -0.39 93.21 +0.06
GCN 92.14 -0.56 89.51 +0.6 98.88 +1.09 94.21 -0.49 91.00 +0.08 95.50 +0.51 91.87 +0.27 93.49 +0.34
GAT 92.79 +0.09 90.49 +1.58 97.90 +0.11 94.03 -0.67 91.76 +0.84 95.13 +0.14 91.92 +0.32 93.53 +0.38
ECL-GSR 93.32 +0.62 84.75 -4.16 97.91 +0.12 94.60 -0.1 92.59 +1.67 95.96 +0.97 91.88 +0.28 92.94 -0.2
MissionHD 93.41 +0.71 91.07 +2.16 98.32 +0.53 94.45 -0.25 91.29 +0.37 95.16 +0.17 91.95 +0.35 93.71 +0.56
@@ 5.1 Experimental Settings
We evaluate our proposed MissionHD framework on the challenging tasks of VAD and VAR with an NVIDIA A100 GPU. Our experimental setup closely follows the methodology established in the previous paper Yun et al. (2025). We conduct experiments on two widely-used benchmarks for real-world anomaly detection: UCF-Crime Sultani et al. (2018) and XD-Violence Wu et al. (2020), both of which consist of videos captured from surveillance cameras. For pre-trained joint embedding model, we used ImageBind-Huge Girdhar et al. (2023). Also, GPT-4 is used to generate LLM-generated MSGs. Performance is measured using the Area Under the Receiver Operating Characteristic Curve (AUC) for individual classes and the mean AUC (mAUC) across all anomaly classes to provide a comprehensive assessment. Key hyperparameters for our model include a hyperdimensional space of $D=10,000$, a learnable dimensionality of $d=32$ for the learnable hypervectors, a hidden dimensionality of 32 for the final decision transformer layer, and a learning rate of $10^{-5}$. We explore graph refinement thresholds of 0.1, 0.2, 0.4, and 0.8. While our primary evaluation focuses on VAD/VAR, the MissionHD framework is designed to be task-agnostic and can be readily applied to other domains requiring structured, interpretable reasoning over learned knowledge graphs.